# The $\alpha$–regression for compositional data: a unified framework for standard, spatially-lagged, spatial autoregressive and geographically-weighted regression models

### v2

@@ 2.3 Extension of the $\alpha$–regression and related models
. The $\alpha$–regression is based upon minimization of the $L_{2}$ norm (7). In a similar fashion we can minimize the $L_{1}$ norm, yielding the $\alpha$–minium absolute deviations ($\alpha$–MAD) regression
$\text{MAD}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{1}=\sum_{i=1}^{n}\sum_{j=1}^{d}|\mathbf{y}_{ij,\alpha}-\mathbf{\mu}_{ij,\alpha}|.$ (14)
@@
Following tsagris2025 we formulated the $\alpha$–MAD regression as a univariate regression problem, by using the vectorization operation for the responses and by constructing the design matrix in a suitable manner. To make the estimation efficient we utilized the command nlrq() from the R package quantreg (quantreg2025). This approach exhibits dependence on initialization and does not ensure convergence while use of the R built in functions optim or nlminb is computationally more expensive. For these reasons we decided to not pursue this approach further.
@@ 3.1 The $\alpha$–SLX model
The $\alpha$–SLX model extends the standard $\alpha$-regression by incorporating spatial spillover effects through the covariates. The fitted compositional values are given by:
@@ 3.1.1 The contiguity matrix
The Euclidean distance between any two pairs of latitude and longitude, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$. As mentioned earlier, the locations are first mapped from their polar to their Cartesian coordinates (after transforming the degrees into radians)
@@
The $(i,j)$ elemets of the contiguity matrix $\mathbf{W}$ are then defined as $w_{ij}=\tilde{w}_{ij}/\sum_{j=1}^{n}\tilde{w}_{ij}$.
@@ 3.1.2 Choosing $\alpha$ and $k$
The R package blockCV (blockcv2019) implements spatial cross-validation techniques designed to address the spatial autocorrelation inherent in geographical data. Unlike the traditional 10-fold CV, which can lead to overly optimistic model performance estimates when data points are spatially clustered (roberts2017), the spatial version partitions data into spatially separated training and testing folds. This ensures that the testing data are spatially independent from the training data, providing more realistic assessments of model generalization to new geographic areas.
@@ 3.1.3 Spatial MEs
The spatial MEs (SMEs) consist of three components, the direct, the indirect and the total. The direct MEs measure the impact of a change in the local covariate $x_{k}$ on the local composition component $\mu_{i}$. The following formulas are identical to the standard $\alpha$–regression MEs (10), as they depend only on the $\mathbf{\beta}$ coefficients and do not involve spatial terms.
@@ 3.1.4 Properties of the SMEs
  • The standard error of the SMEs can be computed in a manner similar to the MEs of the $\alpha$–regression and the formulas of their Jacobians can be found in Appendix D.
@@ 3.2.3 Prediction of new values
3.2.4 A non-linear spatial autocorrelation test
Moran’s I test (moran1950) is the oldest and most widespread test of spatial autocorrelation. In the univariate case, the test statistic is basically the Pearson correlation between the response $\mathbf{y}$ and the spatially weighted average $\mathbf{Wy}$. The p-value for the hypothesis of no spatial correlation can be computed asymptotically or via permutations.
@@ 3.3.1 Computing $d_{ij}$ in the weighting scheme
Some researchers tend to compute the Euclidean distance between two pairs of latitude and longitude, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$, $d_{ij}=\sqrt{\left(\nu_{i}-\nu_{j}\right)^{2}+\left(v_{i}-v_{j}\right)^{2}}$. There is a fundamental flaw with this approach which is highlighted by mardia2000, pg. 13. Consider for instance the case of two coordinates whose latitude (or longitude) values are $359^{\circ}$ and $1^{\circ}$. Using the previous naive approach yields a distance between the two values $359^{\circ}-1^{\circ}=358^{\circ}$, but the actual distance between them is only $2^{\circ}$. To account for this, the pair of coordinates must first be transformed into their Euclidean coordinates, prior to the application of the Euclidean distance.
@@ 3.3.4 SMEs
$\frac{\partial\mu_{1}\left(\nu_{i},v_{i}\right)}{\partial x_{k}}=\begin{cases}-\mu_{1}\left(\nu_{i},v_{i}\right)\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)&\text{for }i=1\\ \mu_{\ell}\left(\nu_{i},v_{i}\right)\left[\beta_{i-1,k}\left(\nu_{i},v_{i}\right)-\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)\right]&\text{for }\ell=2,\ldots,D.\end{cases}$ (28)
@@ 4.1 Agricultural economics data set
Data regarding crop productivity in the Greek NUTS II region of Thessaly during the 2017-218 cropping year were supplied by the Greek Ministry of Agriculture, also known as farm accountancy data network (FADN) data. The data refer to a sample of farms and initially they consisted of 20 crops, but after grouping and aggregation they were narrowed down to 5 crops [^6] A larger version of this data set was used in mattas2025. Following the EU Regulation No1166/2008 that establishes a framework for European statistics at the level of agricultural holdings the aggregation took place across different output of crops.. These crops are Cereals, Cotton, Tree crops, Other annual crops and pasture and Grapes and wine. For each of the 168 farms with unique coordinates, the cultivated area in each of these 5 grouped crops is known. Note that due to the existence of zeros, the LRA approach, i.e. the family of the $\alpha$–regression models presented earlier with $\alpha=0$ is not applicable.
Figure 1(a) shows the location of Thessaly region in Greece, and Figure 1(b) shows the locations of the farms. The majority of the farms cultivate cereals and only few farms hold grapes and wine. Specifically, 84.52% of the farms cultivate cereals, 50.00% cultivate Cotton, 40.48% maintain tree crops, 81.55% hold other annual crops and pasture, and finally only 16.67% of the farms own grapes and wine.
Refer to caption Refer to caption
(a) Region of Thessaly within Greece. (b) The locations of the 168 farms.
Figure 1: The Thessaly region in Greece.
@@
Table 1: Agricultural economics data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the three models.
@@
Table 2: Agricultural economics data set: Pearson correlations between each pair of the observed and fitted components for each of the three regression models.
@@ 4.2 Meuse river data set
Table 3: Meuse river data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the three models.
@@
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.206 with a standard error equal to 0.164. Table 4 contains the correlations between each pair of the observed and fitted components for each of the three regression models. The $\alpha$–regression, $\alpha$-SLX and the $\alpha$–SAR model have performed equally well.
Table 4: Meuse river data set: Pearson correlations between each pair of the observed and fitted components for each of the three regression models.
@@ 5 Conclusions
The relevant functions to perform all the $\alpha$–regression models are available in the R package CompositionalSR (compositionalsr2025), which imports the package minpack.lm.
Future research could explore nonparametric spatially varying models for compositional data, as well as hybrid approaches that blend GWR with machine learning techniques for complex compositional systems.
Multivariate extensions of Moran’s I test (lin2023) have been proposed, but perhaps we could generalise the test both in arbitrary dimensions and beyond the linear association. One could substitute the Pearson correlation by the distance correlation szekely2007. Due to the constrained nature of the compositional data, the test should be applied to the residuals of a $\alpha$ type regression model and test whether the residuals and their spatial neighboring averaged residuals are dependent or not. The p-value of independence between $\mathbf{Y}_{\alpha}$ and $\mathbf{WY}_{\alpha}$ can be computed either via permutations (szekely2007) or asymptotically via the $\chi^{2}$ distribution (shen2022).
@@ A.3 The GWR model
GWR has become a widely used technique in spatial statistics for modeling spatially varying relationships. Traditional regression assumes stationarity of relationships across space, but GWR relaxes this assumption by allowing coefficients to vary geographically (brunsdon1996). Meanwhile, compositional data–data sets where variables represent proportions of a whole and are constrained to sum to unity–have gained attention in many disciplines, including environmental sciences, geology, and social sciences. When spatial heterogeneity and compositional constraints intersect, specialized methodological developments are required. The foundational work of fotheringham2002 formalized GWR as a local regression technique that incorporates spatial weighting functions to account for the geographical location of observations.
@@
Appendix B Gradient vector and Hessian matrix for the $\alpha$–regression

### v3

@@ 2.4 Robust extensions of the $\alpha$–regression
. The $\alpha$–regression is based upon minimization of the $L_{2}$ norm (7). In a similar fashion one may choose to minimize the sum of the absolute deviations, yielding the $\alpha$–minimum absolute deviations ($\alpha$–MAD) regression
$\text{MAD}\left(\mathbf{Y},\mathbf{X};\alpha,\mathbf{B}\right)=\sum_{i=1}^{n}\|\mathbf{y}_{i,\alpha}-\mathbf{\mu}_{i,\alpha}\|_{1}=\sum_{i=1}^{n}\sum_{j=1}^{d}|\mathbf{y}_{ij,\alpha}-\mathbf{\mu}_{ij,\alpha}|.$
@@
Following Tsagris (2025b) the $\alpha$–MAD regression was formulated as a univariate regression problem, by using the vectorization operation for the responses and by constructing the design matrix in a suitable manner. To make the estimation efficient the command nlrq() from the R package quantreg (Koenker and others, 2024) was utilized. This approach exhibits dependence on initialization and does not ensure convergence. Alternative models include alternative loss functions. Instead of the $L_{1}$ and $L_{2}$ norms, one may use Tukey’s biweight loss function (Tukey, 1960), Hampel’s loss function (Hampel, 1974) or Barron’s general loss function (Barron, 2019), all available in the R package gslnls (Chau, 2025). A practical limitation of these approaches is the increased computational cost.
2.5 The $\alpha$–regression with continuous and compositional predictors
For convenience purposes, and without loss of generality, we will consider the case of a single composition, denote by $\bm{Z}$. The composition $\bm{Z}$ is first transformed using the $\alpha$–transformation (3), hence denoted by $\bm{Z}_{\alpha}$. Principal component analysis computes the eigenvectors $\bm{V}_{\alpha}$ of $\bm{Z}_{\alpha}$ and then the projections on to this orthonormal basis are computed, $\bm{S}_{\alpha}=\bm{Z}_{\alpha}\bm{V}_{\alpha}$. The extension of the $\alpha$–regression to account for compositional predictors is straightforward.
The fitted values are given by
$\mu_{i,\alpha^{\prime}}=\begin{cases}\displaystyle\frac{1}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=1\\[12.91663pt] \displaystyle\frac{e^{\bm{x}^{\top}\bm{\beta}_{i}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{i}}}{1+\sum_{j=1}^{D}e^{\bm{x}^{\top}\bm{\beta}_{j}+\bm{s}_{\alpha^{\prime}}^{\top}\bm{\gamma}_{j}}}&\text{for }i=2,\ldots,D.\end{cases}$ (15)
The notation $\alpha^{\prime}$ highlights that the value of $\alpha$ in the compositional predictors need not be the same as the one used when computing the SSE (7), and to clarify the difference, the new SSE may be written as
$\text{SSE}\left(\bm{Y},\bm{X};\alpha,\alpha^{\prime},\bm{B}\right)=\sum_{i=1}^{n}\|\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\|_{2}=\sum_{i=1}^{n}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\right)^{\top}\left(\bm{y}_{i,\alpha}-\bm{\mu}_{i,\alpha^{\prime}}\right).$ (16)
Unlike the $\alpha$–regression, the extension contains two independent $\alpha$ values, whose values can be chosen via the KLD as mentioned earlier.
@@ 2.5 The $\alpha$–regression with continuous and compositional predictors
In the following sections we define three spatial extensions of the $\alpha$–regression. The models presented cover the Euclidean predictors case, and the inclusion of compositional predictors is straightforward, and hence not covered.
@@ 3.1 The $\alpha$–SLX model
The $\alpha$–SLX model extends the standard $\alpha$–regression by incorporating spatial spillover effects through the covariates. The fitted compositional values are given by:
@@ 3.1.1 The contiguity matrix
As mentioned earlier, the locations (latitude and longitude) between a pair of observations, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$ are first mapped from their polar to their Cartesian coordinates (after transforming the degrees into radians)
@@
The $(i,j)$ elements of the contiguity matrix $\mathbf{W}$ are then defined as $w_{ij}=\tilde{w}_{ij}/\sum_{j=1}^{n}\tilde{w}_{ij}$.
@@ 3.1.2 Choosing $\alpha$ and $k$
The R package blockCV (Valavi et al., 2019) implements spatial cross-validation techniques designed to address the spatial autocorrelation inherent in geographical data. Unlike the traditional 10-fold CV, which can lead to overly optimistic model performance estimates when data points are spatially clustered (Roberts et al., 2017), the spatial version partitions data into spatially separated training and testing folds. This ensures that the testing data are spatially independent from the training data, providing more realistic assessments of model generalization to new geographic areas.
@@ 3.1.3 Spatial MEs
The spatial MEs (SMEs) consist of three components, the direct, the indirect and the total MEs. The following formulas are identical to the standard $\alpha$–regression MEs (10), as they depend only on the $\mathbf{\beta}$ coefficients and do not involve spatial terms.
@@ 3.1.4 Properties of the SMEs
  • The standard error of the SMEs can be computed in a manner similar to the MEs of the $\alpha$–regression and the formulas of their Jacobians can be found in Appendix F.
@@ 3.3.1 Computing $d_{ij}$ in the weighting scheme
Some researchers tend to compute the Euclidean distance between two pairs of latitude and longitude, $(\nu_{i},v_{i})$ and $(\nu_{j},v_{j})$, $d_{ij}=\sqrt{\left(\nu_{i}-\nu_{j}\right)^{2}+\left(v_{i}-v_{j}\right)^{2}}$. There is a fundamental flaw with this approach which is highlighted by Mardia and Jupp, 2000, pg. 13. Consider for instance the case of two coordinates whose latitude (or longitude) values are $359^{\circ}$ and $1^{\circ}$. Using the previous naive approach yields a distance between the two values $359^{\circ}-1^{\circ}=358^{\circ}$, but the actual distance between them is only $2^{\circ}$. To account for this, the pair of coordinates must first be transformed into their Euclidean coordinates, prior to the application of the Euclidean distance.
@@ 3.3.4 SMEs
$\frac{\partial\mu\left(\nu_{i},v_{i}\right)}{\partial x_{k}}=\begin{cases}-\mu_{1}\left(\nu_{i},v_{i}\right)\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)&\text{for }i=1\\ \mu_{\ell}\left(\nu_{i},v_{i}\right)\left[\beta_{i-1,k}\left(\nu_{i},v_{i}\right)-\sum_{j=1}^{d}\beta_{jk}\left(\nu_{i},v_{i}\right)\mu_{j+1}\left(\nu_{i},v_{i}\right)\right]&\text{for }\ell=2,\ldots,D.\end{cases}$ (30)
@@ 4.1 Agricultural economics data set
Data regarding crop productivity in the Greek NUTS II region of Thessaly during the 2017-218 cropping year were supplied by the Greek Ministry of Agriculture, also known as farm accountancy data network (FADN) data. The data refer to a sample of farms and initially they consisted of 20 crops, but after grouping and aggregation they were narrowed down to 5 crops [^6] A larger version of this data set was used in Mattas et al. (2025). Following the EU Regulation No1166/2008 that establishes a framework for European statistics at the level of agricultural holdings the aggregation took place across different output of crops.. These crops are Cereals, Cotton, Tree crops, Other annual crops and pasture and Grapes and wine. For each of the 168 farms with unique coordinates, the cultivated area in each of these 5 grouped crops is known. Note that due to the existence of zeros, the LRA approach, i.e. the family of the $\alpha$–regression models presented earlier with $\alpha=0$ is not applicable.
Figure LABEL:fig_real(a) shows the location of Thessaly region in Greece, and Figure LABEL:fig_real(b) shows the locations of the farms. The majority of the farms cultivate cereals and only few farms hold grapes and wine. Specifically, 84.52% of the farms cultivate cereals, 50.00% cultivate Cotton, 40.48% maintain tree crops, 81.55% hold other annual crops and pasture, and finally only 16.67% of the farms own grapes and wine.
@@
Table 1: Agricultural economics data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the four models.
@@
Table 2: Agricultural economics data set: Pearson correlations between each pair of the observed and fitted components for each of the four regression models.
@@ 4.2 Meuse river data set
Table 3: Meuse river data set: average results regarding the optimal choice of $\alpha$, $k$, $h$, KLD and running time (in seconds) for each of the four models.
@@
We then fitted the regression models using the optimal parameters obtained based on the CV protocol and computed the correlations (component-wise) between the observed and fitted compositions. Table 2 contains these correlations. Evidently, the $\alpha$–SAR model seems to have achieved the best fit. It is important to highlight that the spatial autoregressive parameter $\rho$ of the $\alpha$–SAR model was equal to -0.206 with a standard error equal to 0.164. Table 4 contains the correlations between each pair of the observed and fitted components for each of the four regression models. The $\alpha$–regression, $\alpha$–SLX and the $\alpha$–SAR model have performed equally well.
Table 4: Meuse river data set: Pearson correlations between each pair of the observed and fitted components for each of the four regression models.
@@ 5 Conclusions
The $\alpha$–regression (Tsagris, 2015b) was revisited and its theoretical properties were examined. A robustified version based on the quantile regression was proposed, but it was computationally expensive and convergence issues may arise. The inclusion of compositional predictors was made possible via PCA, but it makes the regression computationally more challenging as it involves a second $\alpha$ parameter.
@@
The R package CompositionalSR (Tsagris, 2025a) was developed to perform all the $\alpha$–regression modes, available to download from CRAN.
Future research could explore nonparametric spatially varying models for compositional data, as well as hybrid approaches that blend GWR with machine learning techniques for complex compositional systems. Further, a spatial non-linear autocorrelation test similar in the spirit of Moran’s I test (Moran, 1950) would be of interest.
@@ A.3 The GWR model
GWR has become a widely used technique in spatial statistics for modeling spatially varying relationships. Traditional regression assumes stationarity of relationships across space, but GWR relaxes this assumption by allowing coefficients to vary geographically (Brunsdon et al., 1996). Meanwhile, compositional data–data sets where variables represent proportions of a whole and are constrained to sum to unity–have gained attention in many disciplines, including environmental sciences, geology, and social sciences. When spatial heterogeneity and compositional constraints intersect, specialized methodological developments are required. The foundational work of Fotheringham et al. (2002) formalized GWR as a local regression technique that incorporates spatial weighting functions to account for the geographical location of observations.
@@
Appendix B Proof of Theorem 2.5
Proof.
The proof proceeds in three steps.
Step 1: Uniform Convergence.
Define the sample objective function
$Q_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\|^{2}\,.$
Using the inequality $|a-b|^{2}\leq 2(|a|^{2}+|b|^{2})$ and Assumption 2.4,
$\mathbb{E}[\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\|^{2}]\leq 2\Big(\mathbb{E}[\|y_{i,\alpha}\|^{2}]+\mathbb{E}\big[\sup_{\theta\in\Theta}\|\mu_{i,\alpha}(\theta)\|^{2}\big]\Big)<\infty\,.$
Thus, $\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\|^{2}$ is integrable and by the Strong Law of Large Numbers (SLLN), for each fixed $\theta\in\Theta$, we have $Q_{n}(\theta)\xrightarrow{a.s.}Q(\theta)\,$ (i.e., point-wise). Since $\Theta$ is compact and $\|y_{i,\alpha}-\mu_{i,\alpha}(\theta)\|^{2}$ is continuous in $\theta$ (as a composition of continuous functions), and dominated by an integrable envelope as shown above, the Uniform Strong Law of Large Numbers (USLLN) implies
$\sup_{\theta\in\Theta}|Q_{n}(\theta)-Q(\theta)|\xrightarrow{a.s.}0.$
Step 2: Well-Separated Minimum.
By Assumption 2.3 and the compactness of $\Theta$, the minimum is well-separated. For any $\epsilon>0$, there exists $\delta>0$ such that:
$\inf_{\theta\in\Theta:\|\theta-\theta_{0}\|\geq\epsilon}Q(\theta)\geq Q(\theta_{0})+\delta.$ (A.2)
Step 3: Consistency Argument.
By definition, the estimator $\hat{\theta}_{n}=\arg\min_{\theta}Q_{n}(\theta)$ satisfies:
$Q_{n}(\hat{\theta}_{n})\leq Q_{n}(\theta_{0}).$ (A.3)
For sufficiently large $n$ and $\delta>0$, Step 1 implies:
$Q_{n}(\theta_{0})\leq Q(\theta_{0})+\frac{\delta}{2}\quad\quad\text{as well as}\quad\quad|Q_{n}(\hat{\theta}_{n})-Q(\hat{\theta}_{n})|<\frac{\delta}{2}\,.$ (A.4)
Suppose, for contradiction, that $\left\|\hat{\theta}_{n}-\theta_{0}\right\|\geq\epsilon$ with positive probability. Then, for sufficiently large $n$:
$Q_{n}(\hat{\theta}_{n})$ $\geq Q(\hat{\theta}_{n})-|Q_{n}(\hat{\theta}_{n})-Q(\hat{\theta}_{n})|$
[…]
$\geq\inf_{\left\|\theta-\theta_{0}\right\|\geq\epsilon}Q(\theta)-\frac{\delta}{2}\quad\text{(by assumption)}$
[…]
which contradicts (A.3). Therefore,
$\mathbb{P}(\|\hat{\theta}_{n}-\theta_{0}\|<\epsilon)\to 1\quad\text{for all }\epsilon>0,$
which establishes $\hat{\theta}_{n}\xrightarrow{p}\theta_{0}$. ∎
Appendix C Proof of Theorem 2.13
Based on the regularity conditions established previously[cite: 11, 13, 17, 18], we now prove that the estimator $\hat{\theta}_{n}$ is asymptotically normally distributed.
Theorem C.1 (Asymptotic Normality).
Under Assumptions 1–7[cite: 11, 12, 13, 17, 18, 19], as $n\to\infty$:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})\xrightarrow{d}N(0,G(\theta_{0})^{-1}\Omega(\theta_{0})G(\theta_{0})^{-1})$ (A.5)
where $G(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$ [cite: 18] and $\Omega(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})r_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$.
Proof.
The proof relies on a Taylor expansion of the first-order condition (FOC) of the objective function $Q_{n}(\theta)$[cite: 9, 10].
Step 1: First-Order Condition.
By Theorem 1, $\hat{\theta}_{n}\xrightarrow{p}\theta_{0}$[cite: 21]. Since $\theta_{0}\in\text{int}(\Theta)$[cite: 11], for sufficiently large $n$, the estimator $\hat{\theta}_{n}$ must satisfy the FOC:
$\nabla Q_{n}(\hat{\theta}_{n})=\frac{2}{n}\sum_{i=1}^{n}g_{i}(\hat{\theta}_{n})^{\top}r_{i}(\hat{\theta}_{n})=0$ (A.6)
where $r_{i}(\theta)=y_{i,\alpha}-\mu_{i,\alpha}(\theta)$ [cite: 9] and $g_{i}(\theta)=-\frac{\partial\mu_{i,\alpha}(\theta)}{\partial\theta}$[cite: 15].
Step 2: Taylor Expansion.
Applying the Mean Value Theorem to $\nabla Q_{n}(\hat{\theta}_{n})$ around the true parameter $\theta_{0}$:
$0=\nabla Q_{n}(\theta_{0})+\nabla^{2}Q_{n}(\bar{\theta})(\hat{\theta}_{n}-\theta_{0})$ (A.7)
where $\bar{\theta}$ lies on the line segment connecting $\hat{\theta}_{n}$ and $\theta_{0}$. Rearranging to solve for the scaled difference:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})=-[\nabla^{2}Q_{n}(\bar{\theta})]^{-1}\sqrt{n}\nabla Q_{n}(\theta_{0})$ (A.8)
Step 3: Convergence of the Hessian (The Bread).
The Hessian of the objective function is given by:
$\nabla^{2}Q_{n}(\theta)=\frac{2}{n}\sum_{i=1}^{n}\left(g_{i}(\theta)^{\top}g_{i}(\theta)+\nabla g_{i}(\theta)^{\top}r_{i}(\theta)\right)$ (A.9)
As $n\to\infty$, $\bar{\theta}\xrightarrow{p}\theta_{0}$. By the Uniform Law of Large Numbers and the fact that $\mathbb{E}[r_{i}(\theta_{0})\mid x_{i}]=0$[cite: 19, 20]:
$\nabla^{2}Q_{n}(\bar{\theta})\xrightarrow{p}2\mathbb{E}[g_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]=2G(\theta_{0})$ (A.10)
where $G(\theta_{0})$ is positive definite[cite: 18].
Step 4: Convergence of the Gradient (The Meat).
The term $\sqrt{n}\nabla Q_{n}(\theta_{0})$ is a sum of i.i.d. random variables[cite: 19]:
$\sqrt{n}\nabla Q_{n}(\theta_{0})=\frac{2}{\sqrt{n}}\sum_{i=1}^{n}g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})$ (A.11)
By the Central Limit Theorem, since $\mathbb{E}[g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})]=0$[cite: 20]:
$\sqrt{n}\nabla Q_{n}(\theta_{0})\xrightarrow{d}N(0,4\Omega(\theta_{0}))$ (A.12)
where $\Omega(\theta_{0})=\mathbb{E}[g_{i}(\theta_{0})^{\top}r_{i}(\theta_{0})r_{i}(\theta_{0})^{\top}g_{i}(\theta_{0})]$.
Step 5: Final Result via Slutsky’s Theorem.
Combining the results using Slutsky’s Theorem:
$\sqrt{n}(\hat{\theta}_{n}-\theta_{0})$ $\xrightarrow{d}-[2G(\theta_{0})]^{-1}N(0,4\Omega(\theta_{0}))$
$=N(0,[2G(\theta_{0})]^{-1}(4\Omega(\theta_{0}))[2G(\theta_{0})]^{-1})$
$=N(0,G(\theta_{0})^{-1}\Omega(\theta_{0})G(\theta_{0})^{-1})$
This completes the proof of asymptotic normality with the sandwich covariance matrix. ∎
Appendix D Gradient vector and Hessian matrix for the $\alpha$–regression